Which of the following gives the equation that relates the rate of consumption of Al to the rate of formation of Al2O3 for the reaction below?

8Al+3Fe3O4→4Al2O3+9Fe

Select the correct answer below:

1/8Δ[Al]Δt=14Δ[Al2O3]Δt

−1/8Δ[Al]Δt=14Δ[Al2O3]Δt

−8Δ[Al]Δt=4Δ[Al2O3]Δt

8Δ[Al]Δt=4Δ[Al2O3]Δt
Answer #1

rate of reaction is given as:

rate = sign * (1/x) * d[concentration]/dt

sign is negative for reactant and positive for product

x is stoichiometric coefficient in chemical equation

Rate in terms of Al is:

Rate = (-1/8)Δ[Al]/Δt

Rate in terms of Al2O3 is:

Rate = (1/4)Δ[Al2O3]/Δt

Equate both:

(-1/8)Δ[Al]/Δt = (1/4)Δ[Al2O3]/Δt

Answer: option 2